OpenStudy (anonymous):

Write an equation and solve. Two times the difference of a number decreased by five is fourteen. What is the number?

OpenStudy (anonymous):

2x-5=14

OpenStudy (anonymous):

Solve for x: 2 x-5 = 14 Isolate terms with x to the left hand side. Add 5 to both sides: 2 x+(5-5) = 14+5 Look for two terms that sum to zero. 5-5 = 0: 2 x = 14+5 Evaluate 14+5. 14+5 = 19: 2 x = 19 Solve for x. Divide both sides of 2 x = 19 by 2: Answer: x = 19/2

OpenStudy (anonymous):

hmm u sure ?

OpenStudy (anonymous):

Isn't this saying 2(x -5)=14 ?

OpenStudy (epikrebel):

I was going to say that.

OpenStudy (epikrebel):

So x=12

OpenStudy (anonymous):

haha, oh I didn't even look at the original question that hard just took her equation

OpenStudy (anonymous):

Solve for x: 2 (x-5) = 14 Divide both sides by a constant to simplify the equation. Divide both sides of 2 (x-5) = 14 by 2: x-5 = 14/2 Divide 14 by 2. 14/2 = (2×7)/2 = 7: x-5 = 7 Isolate terms with x to the left hand side. Add 5 to both sides: x+(5-5) = 7+5 Look for two terms that sum to zero. 5-5 = 0: x = 7+5 Evaluate 7+5. 7+5 = 12: Answer: x = 12

OpenStudy (anonymous):

Write an equation and solve. Five times the difference of a number subtracted from ten is forty. What is the number?

OpenStudy (anonymous):

The question was confusingly worded IMHO

OpenStudy (anonymous):

Solve for x: 5 (x-10) = 40 Divide both sides by a constant to simplify the equation. Divide both sides of 5 (x-10) = 40 by 5: x-10 = 40/5 Divide 40 by 5. 40/5 = (5×8)/5 = 8: x-10 = 8 Isolate terms with x to the left hand side. Add 10 to both sides: x+(10-10) = 8+10 Look for two terms that sum to zero. 10-10 = 0: x = 8+10 Evaluate 8+10. 8+10 = 18: Answer: x = 18

OpenStudy (epikrebel):

x=18